# LLMs hallucinate graphs too: a structural perspective

### v1

@@ Abstract
It is known that LLMs do hallucinate, that is, they return incorrect information as facts. In this paper, we introduce the possibility to study these hallucinations under a structured form: graphs. Hallucinations in this context are incorrect outputs when prompted for well known graphs from the literature (e.g. Karate club, Les Misérables, graph atlas). These hallucinated graphs have the advantage of being much richer than the factual accuracy –or not– of a fact; this paper thus argues that such rich hallucinations can be used to characterize the outputs of LLMs. Our first contribution observes the diversity of topological hallucinations from major modern LLMs. Our second contribution is the proposal of a metric for the amplitude of such hallucinations: the Graph Atlas Distance, that is the average graph edit distance from several graphs in the graph atlas set. We compare this metric to the Hallucination Leaderboard, a hallucination rank that leverages $10,000$ times more prompts to obtain its ranking.
@@ 1 Introduction
While querying LLMs for graph structures is unexplored and might sound promising, it raises a number of questions we further investigate in the remainder of this paper. First, we detail the envisioned interaction framework in which the objective is to collect a graph as a LLM output. Once this basic interaction framework has been established, a first research question refers to the quality of the collected graphs. Namely, are output graphs accurate with regard to the original queried graph? And moreover, are some answers wrong and inaccurate, in other words do LLMs hallucinate graphs? Second, we showcase the potential of graph requests and their capacity to extract information regarding target LLMs in a request-efficient way. To that end, we compare our proposal (the Graph Atlas Distance) to the Hallucination Leaderboard [4].
@@ 2.1 Statistics on the Topology of Output Graphs by LLMs
Finally we report in the last column the $l2$ distance between the degree distribution of the output graph and KC. Please note that except for rare unigraphic graphs, there exists multiple graphs having the same degree distribution, so this metric is not a correctness assessment per se. In other words, while strictly positive such distance implies a incorrect result, a distance of 0 does not imply a perfect (isomorphic to KC) result. The table is sorted using this last column.
Table 1 shows that no output graph is exactly correct: all LLMs do hallucinate. The closest one being by the dbrx-instruct model, with just two hallucinated (i.e. added) edges. ChatGPT models (3.5 and 4o) then follow and both return the same output graph. Many results are radically different from the correct one (hence deserving being coined hallucinations), with a surprisingly high heterogeneity among the answers (for instance, with $|E|$ ranging from 8 to 153).
@@
The bottom of the list is populated by output graphs that heavily differ from the ground truth, with i.e. the graph on Figure 2a that is a star with two central nodes (resulting in a low assortativity score). In this precise case, the number of nodes is correct, 34, and the number of edges close to correct, showing that model has probably learned these global features but not the topology. Figure 2b demonstrate the opposite case for llama-2-13b-chat-hf, where node and edge counts are very low (6 and 8 respectively). Another interesting case is mistral-large having the correct number of nodes while hallucinating nearly twice the number of edges (see Figure 2c), explaining its high distance to the degree sequence of KC, and its poor modularity in comparison. Finally, Figure 2d demonstrates for snowflake-arctic-instruct the output of a dense graph (except for a node, resulting in a high density).
On the model family side, we observe with llama variants that a higher amount of parameters does not imply more correctness: lama-3.1-405B-Instruct-Turbo is ranked in the middle in that regard. We also observe that for a model version (llama-3-70B), different variants (lite, turbo and groq) result in related ranks, yet with significantly different characteristic outputs.
@@ 2.2 Embedding LLMs Output Graph and the Ground Truth
We now rely on graph embeddings, for examining the relative proximity of LLM outputs. We employ the Karate Club library [6] for that purpose, with the NetLSD method [21] with default parameters, before we run a t-SNE appearing in Figure 3. (Les Misérables and graph atlas 50 are deferred in Appendix B.)
@@
Despite the proximity of some related model architectures such as llama-2 or ChatGPT, a relatively smooth spread appear in this t-SNE, with clusters that mix graphs obtained by different vendors. Older versions of llama models (version 2) appear the less related to KC in this embedding.
@@ 3 A LLM Hallucination Rank Based on Output Graphs
Benchmarks for measuring hallucinations are constituted of tens of thousand of prompts [4, 17]. They hence require a privileged access to the model, compared to web browser based prompting access. We here propose to sequentially prompt a LLM for just a handful of graphs from the graph atlas, and to average errors made regarding this reference under a single value (using the average error of graph edit distances).
@@
We note that across all tested LLMs and for all datasets, an isomorphism solely occurred for gpt4o on graph atlas #7 (a triangle graph) and #13 (a star composed of 4 nodes). Regarding the weight we give to operation in the graph edit distance, we do not account for labels, but consider node/edge insertion/deletion each costing 1.
We compare against the Hallucination Leaderboard [4], a GitHub page ranking LLMs based on the amplitude of their hallucinations, using a dataset of 50k prompts. The rank of the ten tested graphs in common with the Hallucination Leaderboard are presented in Table 4.
@@
Figure 4: GAD (resolution: 5 graphs) vs a summarization of a benchmark based on prompts with binary answers [4]. In the GAD column, the first number is the average edit distance distance, the second is the standard deviation. ($\star$ rank in [4] as been adjusted to the LLMs available in our experiment)
We can observe an interesting correlation in these two rankings (with a spearman rank correlation of $0.3$, where 00 is random). The first position is held by gpt4o. The first of the 4 llama models (llama-3.1-70B-Instruct) is down-ranked in GAD, yet the 3 others are in the correct order. The larger 405B parameters llama perform worse than smaller models; note that this inversion also appears in [4], where llama 3 beats llama 3.1 with the same amount of parameters (70B). qwen2-72B-Instruct ranks in the middle. c4ai-command-r-plus ranks at the bottom. snowflake-arctic-instruct is nevertheless strongly down-ranked with GAD, as compared to its second position in the Hallucination Leaderboard.
@@ 4 Discussion
Our approach explores graph requests as a mean to obtain more information bits by prompt, compared to the standard MCQ approach naturally in which each request yields a number of bits capturing the number of possible answers to each question (typically 2 or 4 choices, hence 1 or 2 bits).
Implicitly, this approach hence compares interrogation patterns "bit to bit", regardless of the relevance of the collected bits. A natural limitation is to overlook the precise meaning of each bit. Concretely, for instance, one bit yielded by our approach captures whether the target LLM correctly predicted a (Jean Valjean, Cosette) relation in Les Misérables, whereas one bit yielded by MedMCQA benchmarks captures whether the target LLM correctly associates "A 40-year-old man has megaloblastic anemia and early signs of neurological abnormality" to a deficit in B12 Vitamin.
The difficulty with considering the semantics associated with each bit is that the notion of relevance is strongly application dependent. The relevance varies depending on the use case, be it for fingerprinting a target LLM or for evaluating its ability to provide truthful medical advice.
@@ 6 Conclusion
This paper made the case for comparing ground truth graphs from the literature to the output of LLMs, prompted to output them. A first striking observation is that current LLMs are far from being a reliable source of edge lists (i.e. they do hallucinate graphs). Nevertheless, these glitches open solid comparison avenues, which we exploited to observe the significant differences of LLMs in their hallucination amplitude regarding standard graphs. Introducing relevant metrics, we have illustrated that a handful of prompts to a LLM can correlate with a method leveraging tens of thousands of queries for binary answers, in the capability to rank the less hallucinating LLMs. We believe that such a novel perspective leaves ways for future work to constitute robust and more efficient benchmarks for assessing the quality of LLMs.
@@ Appendix C Spectral Distances from Ground Truth Graphs
with $s$ the set of eigenvalues $s=\{\lambda_{1},\lambda_{2},\dots,\lambda_{|V|}\}$, knowing that $\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{|V|}$. As recommended in [22], if graphs are of different sizes, the missing eigenvalues of the smaller are zeros padded.

### v2

@@ Abstract
It is known that LLMs do hallucinate, that is, they return incorrect information as facts. In this paper, we introduce the possibility to study these hallucinations under a structured form: graphs. Hallucinations in this context are incorrect outputs when prompted for well known graphs from the literature (e.g. Karate club, Les Misérables, graph atlas). These hallucinated graphs have the advantage of being much richer than the factual accuracy –or not– of a statement; this paper thus argues that such rich hallucinations can be used to characterize the outputs of LLMs. Our first contribution observes the diversity of topological hallucinations from major modern LLMs. Our second contribution is the proposal of a metric for the amplitude of such hallucinations: the Graph Atlas Distance, that is the average graph edit distance from several graphs in the graph atlas set. We compare this metric to the Hallucination Leaderboard, a hallucination rank that leverages $10,000$ times more prompts to obtain its ranking.
@@ 1 Introduction
While querying LLMs for graph structures is unexplored and might sound promising, it raises a number of questions we further investigate in the remainder of this paper. First, we detail the envisioned interaction framework in which the objective is to collect a graph as a LLM output. Once this basic interaction framework has been established, we first focus on the quality of the collected graphs. Namely, are output graphs accurate with regard to the original queried graph? And moreover, are some answers wrong and inaccurate, in other words do LLMs hallucinate graphs? Second, we showcase the potential of graph requests and their capacity to extract information regarding target LLMs in a request-efficient way. To that end, we compare our proposal (the Graph Atlas Distance) to the Hallucination Leaderboard [4].
@@ 2.1 Statistics on the Topology of Output Graphs by LLMs
Finally we report in the last column the $l2$ distance between the degree distribution of the output graph and KC. Please note that except for rare unigraphic graphs, there exists multiple graphs having the same degree distribution, so this metric is not a correctness assessment per se. In other words, while a positive distance implies an incorrect result, a distance of 0 does not imply a perfect result (i.e. isomorphic to KC). The table is sorted using this last column.
Table 1 shows that no output graph is exactly correct: all LLMs do hallucinate. The closest one being by the dbrx-instruct model, with just two hallucinated (i.e. added) edges. ChatGPT models (3.5 and 4o) then follow and both return the same output graph. Many outputs are radically different from the correct one (hence deserving being coined hallucinations), with a surprisingly high heterogeneity among the answers (for instance, with $|E|$ ranging from 8 to 153).
@@
The bottom of the list is populated by output graphs that heavily differ from the ground truth, with i.e. the graph on Figure 2a that is a star with two central nodes (resulting in a low assortativity score). In this precise case, the number of nodes is correct, 34, and the number of edges close to correct, showing that model has probably learned these global features but not the topology. Figure 2b demonstrates the opposite case for llama-2-13b-chat-hf, where node and edge counts are very low (6 and 8 respectively). Another interesting case is mistral-large having the correct number of nodes while hallucinating nearly twice the number of edges (see Figure 2c), explaining its high distance to the degree sequence of KC, and its poor modularity in comparison. Finally, Figure 2d showcases for snowflake-arctic-instruct the output of a dense graph (except for a node, resulting in a high density).
Regarding model families, we observe with variants of llama that a higher number of parameters does not imply increased correctness. Indeed, the largest llamma, lama-3.1-405B-Instruct-Turbo, is ranked in the middle. We also observe that for a model version (llama-3-70B), different variants (lite, turbo and groq) result in related ranks, yet with significantly different characteristic outputs.
@@ 2.2 Embedding LLMs Output Graph and the Ground Truth
We now rely on graph embeddings, for examining the relative proximity of LLM outputs. We employ the Karate Club library [6] for that purpose, with the NetLSD method [21] with default parameters, before we run a t-SNE appearing in Figure 3. (Les Misérables and graph atlas 50 representations are deferred in Appendix B.)
@@
Despite the proximity of model families such as llama-2 or ChatGPT, a relatively smooth spread appear in this t-SNE, with clusters that mix graphs obtained by different vendors. Older versions of llama models (version 2) appear the less related to KC in this embedding.
@@ 3 A LLM Hallucination Rank Based on Output Graphs
Benchmarks for measuring hallucinations are constituted by tens of thousand of prompts [4, 17]. They hence require a privileged access to the model, compared to web browser based prompting access. We here propose to sequentially prompt a LLM for just a handful of graphs from the graph atlas, and to average errors made regarding this reference under a single value (using the average of graph edit distances).
@@
We note that across all tested LLMs and for all datasets, an isomorphism solely occurred for gpt4o on graph atlas #7 (a triangle graph) and #13 (a star composed of 4 nodes). The weight given to edit operations are is follows: we do not account for labels, but consider node/edge insertion/deletion each costing 1.
We compare against the Hallucination Leaderboard [4], a GitHub page ranking LLMs based on the amplitude of their hallucinations, using a dataset of 50k prompts. The rank of the 10 tested graphs in common with the Hallucination Leaderboard is presented in Table 4.
@@
Figure 4: GAD (resolution: 5 graphs) vs a summarization of a benchmark based on prompts with binary answers [4]. In the GAD column, the first number is the average edit distance distance, the second is the standard deviation. ($\star$ rank in [4] as been adjusted to the LLMs available in our experiment).
We can observe an interesting correlation in these two rankings (with a spearman rank correlation of $0.3$, where 00 is random). The first position is held by gpt4o. The first of the 4 llama models (llama-3.1-70B-Instruct) is down-ranked in GAD, yet the 3 others are in the same order. The larger 405B parameters llama perform worse than smaller models; note that this inversion also appears in [4], where llama 3 beats llama 3.1 with the same amount of parameters (70B). qwen2-72B-Instruct similarly ranks in the middle. c4ai-command-r-plus ranks at the bottom. snowflake-arctic-instruct is nevertheless strongly down-ranked with GAD, as compared to its second position in the Hallucination Leaderboard.
@@ 4 Discussion
Our approach explores graph requests as a mean to obtain more information bits by prompt, compared to the standard MCQ approach in which each request yields a number of bits capturing the number of possible answers to each question (typically 2 or 4 choices, hence 1 or 2 bits). Implicitly, this approach hence compares interrogation patterns "bit to bit", regardless of the relevance of the collected bits. A natural limitation is to overlook the precise meaning of each bit. Concretely, for instance, one bit yielded by our approach captures whether the target LLM correctly predicted a (Jean Valjean, Cosette) relation in Les Misérables, whereas one bit yielded by MedMCQA benchmarks captures whether the target LLM correctly associates "A 40-year-old man has megaloblastic anemia and early signs of neurological abnormality" to a deficit in B12 Vitamin. The difficulty with considering the semantics associated with each bit is that the notion of relevance is strongly application dependent. The relevance varies depending on the use case, be it for fingerprinting a target LLM or for evaluating its ability to provide truthful medical advice.
@@ 6 Conclusion
This paper made the case for comparing ground truth graphs from the literature to the output of LLMs, prompted to output them. A first striking observation is that current LLMs are far from being reliable sources of edge lists (i.e. they do hallucinate graphs). Nevertheless, these glitches open solid comparison avenues, which we exploited to observe the significant differences of LLMs in their hallucination amplitude regarding standard graphs. Introducing relevant metrics, we have illustrated that a handful of prompts to a LLM can correlate with a method leveraging tens of thousands of queries for binary answers, in the capability to rank the less hallucinating LLMs. We believe that such a novel perspective leaves ways for future work to constitute robust and more efficient benchmarks for assessing the quality of LLMs.
@@ Appendix C Spectral Distances from Ground Truth Graphs
with $s$ the set of eigenvalues $s=\{\lambda_{1},\lambda_{2},\dots,\lambda_{|V|}\}$, knowing that $\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{|V|}$. As recommended in [22], if graphs are of different sizes, the missing eigenvalues of the smaller are padded with zeros.